# 3D Gaussian Splatting: Survey, Technologies, Challenges, and Opportunities

### v1

@@ 7.1 Suboptimal Data Challenges
In real-world scenarios, collecting a large volume of high-quality training data is often impractical. On one hand, without access to 3D data and sufficient multi-view images, relying solely on a limited number of 2D image supervisions is insufficient for accurate 3DGS reconstruction. For instance, with only a frontal image of a target, inferring the appearance of the back is highly challenging. On the other hand, data quality is equally critical; accurate poses and clear images directly influence reconstruction performance.
Such problems are discussed across multiple tasks, such as Sparse Views Setting (Sec.3.3.1), Autonomous Driving (Sec.4.3), Dynamic 3DGS (Sec.5.1) (Monocular Video), AIGC (Sec.4.2), and Editable 3DGS (Sec.5.3). Numerous works have discussed how to improve initialization (Sec.6.1), regularization (Sec.6.4), adaptive control (Sec.6.6), and training strategies (Sec.6.5), or introduce additional priors (Sec.6.9) in the context of sparse or missing multi-view training images. Additionally, reconstruction with missing accurate poses can also be considered an underdetermined problem. This is discussed in the SLAM literature (Sec.4.3.2) and reconstruction under blurred images (Sec.3.2) or without poses [207, 208], where new training strategies (Sec. 6.5) and other representations (Sec. 6.8) are often introduced to mitigate it.
7.2 Generalization Challenges
Despite the improved training efficiency compared to NeRF, the scene-specific training paradigm remains a major bottleneck for the application of 3DGS. It is hard to imagine having to train for each target or scene individually, especially in multi-target reconstruction and text-to-scene generation.
We find that this issue is commonly discussed in the Generalization setting (Sec.3.3.1) and in some Generalization-related tasks in Human Reconstruction (Sec.4.1) and AIGC (Sec.4.2). These works aim to directly infer new scenes using feedforward networks. Therefore, they exhibit strong technical similarities, such as the introduction of triplanes. Specific improvement strategies are extensively analyzed in initialization (Sec.6.1), adaptive control (Sec.6.6), and integration with other representations (Sec.6.8).
7.3 Physics Challenges
Traditional 3DGS only considers static rendering and neglects the laws of physics motion, which are important in simulations [139]. Additionally, physically-based rendering is a significant step towards applying 3DGS to simulate the physical world. Thanks to the explicit representation, it is possible to construct 3DGS that adhere to physical laws.
This issue is commonly discussed in Dynamic 3DGS (Sec.5.1), Physics Simulation (Sec.5.5), Animation (Sec.5.3.4), and some dynamic Human Reconstruction (Sec.4.1) or Autonomous Driving Scene Reconstruction (Sec.4.3.1). These discussions focus on how Gaussian primitives should move in the physical world, and physically-based rendering is introduced in Photorealism (Sec.3.2). From a technical perspective, since the early stages of 3DGS research, numerous works have addressed this issue, often focusing on attribute expansion (Sec.6.2), regularization strategy (Sec.6.4), and the assistance of additional priors (Sec. 6.9).
7.4 Realness and Efficiency Challenges
Realness and efficiency challenges are fundamental issues. They are investigated in various works and have been discussed in Sec. 3. In this section, we discuss some typical related tasks and summarize their common techniques.
Regarding the Realness challenges, existing works have discussed not only Photorealism (Sec.3.2), Surface Reconstruction (Sec.5.2), and Semantic Understanding (Sec.5.4), but also addressed this issue in AIGC-related (Sec.4.2) and Autonomous Driving (Sec.4.3) research. Most of the techniques mentioned in (Sec.6) contribute to the improvement of rendering performance. Different methods are employed for different tasks. For instance, AIGC-related works (Sec.4.2) often focus on improving training strategies (Sec.6.5) and regularization (Sec.6.4), while Surface Reconstruction works (Sec.5.2) are related to post-processing procedures (Sec. 6.7).
Regarding the efficiency challenges (Sec.3.1), they are mentioned in some derivative tasks, such as those related to Autonomous Driving (Sec.4.3) and Semantic Understanding (Sec.5.4). Existing improvements often focus on introducing additional attributes (Sec.6.2) or post-processing (Sec.6.7), as well as improving adaptive control (Sec.6.6) and splatting (Sec. 6.3) strategies.
In addition, there are some relationships between different domains that have not been mentioned. For instance, Surface Reconstruction techniques (Sec.5.2) are often referenced in the context of Editable 3DGS (Sec.5.3), etc.
8 Opportunities
3DGS has recently experienced significant development, with numerous works demonstrating its potential in related tasks. In this section, we discuss unresolved issues among the aforementioned core challenges and propose potential directions for future studies. Additionally, we discuss some extensions of applications and technologies.
8.1 Suboptimal Training Data
An ideal 3DGS training process necessitates sufficient and high-quality data, but in practical applications, this is often excessively challenging. Although focusing on the introduction of priors can mitigate the problem to some extent, optimizing a large number of Gaussians under underconstrained conditions remains inherently difficult. Therefore, a potential solution is to reduce the number of Gaussian primitives based on their uncertainty while enhancing the representational capacity of individual primitives [27]. This involves finding a trade-off between the number of Gaussians and rendering performance, thereby improving the efficiency of utilizing sparse samples.
@@ 8.1 Suboptimal Training Data
Then, poor-quality data should also be taken into consideration. Unconstrained in-the-wild images are a typical case, encompassing transient occlusions and dynamic appearance changes, such as varying sky, weather, and lighting, which have been extensively discussed in NeRF [209, 210, 211]. To enhance efficiency, existing works have also discussed this issue in the context of 3DGS [212, 213], which attempt to model appearance changes and handle transient objects. However, their performance struggles, especially in scenes with complex lighting changes and frequent occlusions. Thanks to the explicit representation characteristics of 3DGS, decoupling geometric representations and introducing geometric consistency constraints across different scenes is a promising approach to mitigate instability during the training process.
8.2 Generalization
@@
Although existing generalization-related works can directly obtain scene representations through forward inference, their performance is often unsatisfactory and limited by the type of scene [49, 93, 46, 50]. We hypothesize that this is due to the difficulty of feedforward networks in performing adaptive control of 3DGS, as also mentioned in [47]. In future research, designing a reference-feature-based feedforward adaptive control strategy is a potential solution, which can predict the positions requiring adaptive control through reference features and be plug-and-play into existing generalization-related works. Additionally, existing generalization-related works rely on accurate poses, which are often difficult to obtain in practical applications [214, 208, 215]. Therefore, discussing generalizable 3DGS under pose-missing conditions is also promising [204].
8.3 Physics Problems
8.3.1 Physically-based Motion
@@
Ensuring that the motion of 3DGS adheres to physical laws is essential for unifying simulation and rendering [139]. Although rigidity-related regularization terms have been introduced, as described in Sec.6.4.1, most existing works focus on animating 3DGS while neglecting the physical attributes of the Gaussian primitives themselves (Sec.5.1). Some pioneering works attempt to introduce velocity attributes [106] and Newtonian dynamics rules [139], but this is not sufficient to fully describe the physical motion of 3DGS in space. A potential solution is to introduce more physical attributes in Gaussian primitives, such as material [216], acceleration, and force distribution, which can be constrained by priors from certain simulation tools and physics knowledge.
8.3.2 physically-based rendering
Physically-based rendering is also a direction worth attention, as it enables 3DGS to handle relighting and material editing, producing outstanding inverse rendering results [40]. Future works can explore decoupling geometry and appearance in 3DGS, conducting research from the perspectives of normal reconstruction and the modeling of illumination and materials [216, 217, 72].
8.4 Realness and Efficiency
8.4.1 Surface Reconstruction
@@
The difficulty in reconstructing clear surfaces has always been a significant challenge affecting rendering realism. As discussed in Sec. 5.2, some works have addressed this issue, attempting to represent surfaces with planar Gaussians. However, this can result in a decline in rendering performance, possibly due to the reduced representational capacity of planar Gaussian primitives or the instability of the training process. Therefore, designing Gaussian primitives that are better suited for surface representation and introducing a multi-stage training paradigm along with regularization terms are potential solutions.
8.4.2 Efficiency
Storage efficiency is one of the critical bottlenecks of 3DGS. Existing works focus on introducing VQ techniques and compressing SH parameters, as discussed in Sec. 3.1.1. However, such approaches inevitably affect rendering performance. Therefore, exploring how to design more efficient representations based on 3DGS is a potential way to enhance efficiency [26, 27], while maintaining performance.
8.5 Application
8.5.1 Large-scale Gaussian Splatting
Although recent methods are capable of efficient reconstruction in small-scale and object-centric scenes, extending this capability to large-scale scenes remains challenging due to limited video memory, lengthy optimization times, and significant appearance changes. Some recent works have attempted to address these issues. For example, VastGaussian [218] intuitively uses a region division mechanism based on camera positions to divide large scenes into multiple small blocks and introduces additional auxiliary camera positions based on a spatially-aware visibility criterion, thus achieving precise reconstruction for each small region. Similarly, Fed3DGS [219] introduces a federated learning framework into large-scale 3DGS to reduce the data load on the central server and implements a Gaussian merging strategy, enabling decentralized reconstruction among millions of clients and distributed computing resources. Following these approaches, there remains substantial research space.
8.5.2 AIGC
Existing AIGC-related works mostly focus on the generation of single static objects. However, in practical applications, the generation of multiple objects (Sec.4.2.3) and dynamic objects (Sec.4.2.4) are more important. Additionally, in constructing scenes with multiple moving objects, characterizing the objects interactions is also a worthwhile research task.
8.5.3 Application Expansion
Thanks to the efficient and robust reconstruction capabilities of 3DGS, numerous applications have garnered attention in various fields, such as medicine [220], industrial defect detection [221], image compression [222], and aviation [223]. In the future, 3DGS could potentially replace NeRF for 3D reconstruction in additional fields, such as robotics [224] and biology [225].
8.6 Technical Improvements
8.6.1 Initialization
@@
An increasing number of research and engineering projects have found that initialization is important in 3DGS. Traditional SfM initialization is not suitable for many constrained scenarios, such as sparse view settings, AIGC, and low-light reconstruction. Therefore, more robust initialization methods should be designed to replace random initialization in these constrained scenarios.
8.6.2 Splatting
Splatting also plays an important role in 3DGS, but it is rarely mentioned in existing works [191, 190]. Designing efficient parallel splatting strategies on pretrained 3DGS has the potential to impact rendering performance and efficiency.
9 Conclusion
@@
The burgeoning interest in the field of 3D Gaussian Splatting (3DGS) has precipitated the emergence of a myriad of related downstream tasks and technologies, thereby contributing to an increasing complexity and confusion within the domain, which manifest in various forms, including the similar motivations across different works; the incorporation of similar technologies across disparate tasks; and the nuanced differences and interconnections between various technologies. In this survey, we endeavor to systematically categorize existing works based on their underlying motivations and critically discuss the associated technologies. Our objective is to elucidate common challenges across different tasks and technologies, thereby providing a coherent framework for comprehending this rapidly evolving field. This survey aspires to serve as an invaluable resource for both newcomers and seasoned practitioners, facilitating navigation and effective engagement with recent advancements. Additionally, we identify and highlight prospective avenues for future research, aiming to inspire continued innovation and exploration in 3DGS.

### v2

@@ VIII-B Suboptimal Data
Challenges. In real-world scenarios, collecting large volumes of high-quality training data is often impractical. Without access to 3D data and sufficient multi-view images, relying on limited 2D image supervision is insufficient for accurate 3DGS reconstruction. For example, inferring the back appearance from only a frontal image is highly challenging. Additionally, data quality is critical, as accurate poses and clear images directly influence reconstruction performance.
@@
Opportunities. An ideal 3DGS training process requires sufficient high-quality data, but this is often excessively challenging in practical applications. Although introducing priors can mitigate this problem to some extent, optimizing a large number of Gaussians under underconstrained conditions remains difficult. A potential solution is to reduce the number of Gaussian primitives based on their uncertainty while enhancing the representational capacity of individual primitives [17]. This involves finding a trade-off between the number of Gaussians and rendering performance, thereby improving the efficiency of utilizing sparse samples. Then, poor-quality data should also be taken into consideration. Unconstrained in-the-wild images are a typical case, encompassing transient occlusions and dynamic appearance changes, such as varying sky, weather, and lighting, which have been extensively discussed in NeRF [266, 267, 268]. To enhance efficiency, existing works have addressed this issue in the context of 3DGS [269, 270], attempting to model appearance changes and handle transient objects. However, their performance struggles, especially in scenes with complex lighting changes and frequent occlusions. Thanks to the explicit representation characteristics of 3DGS, decoupling geometric representations and introducing geometric consistency constraints across different scenes is a promising approach to mitigate instability during the training process.
VIII-C Generalization
Challenges. Despite the improved training efficiency compared to NeRF, the scene-specific training paradigm remains a major bottleneck for the application of 3DGS. It is hard to imagine having to train for each target or scene individually, especially in multi-target and scene reconstruction (generation).
@@ VIII-C Generalization
Opportunities. Although existing generalization-related works can directly obtain scene representations through forward inference, their performance is often unsatisfactory and limited by the type of scene [47, 114, 45, 51]. We hypothesize that this is due to the difficulty of feedforward networks in performing the adaptive control of 3DGS, as also mentioned in [46]. In future research, designing a reference-feature-based feedforward adaptive control strategy is a potential solution, which can predict the positions requiring adaptive control through reference features and be plug-and-play into existing generalization-related works. Additionally, existing generalization-related works rely on accurate poses, which are often difficult to obtain in practical applications [271, 265, 272]. Therefore, discussing generalizable 3DGS under pose-missing conditions is also promising [259].
VIII-D Physics Reconstruction and Rendering
Challenges. Traditional 3DGS only considers static rendering and neglects the laws of physical motion, which are important in simulations [191]. Additionally, Physically-based rendering is a significant step towards applying 3DGS to simulate the physical world and achieve more realistic effects.
@@ VIII-D Physics Reconstruction and Rendering
Opportunities. Ensuring that the 3DGS’s motion adheres to physical laws is essential for unifying simulation and rendering [191]. Although rigidity-related regularization have been introduced, as described in Sec. VI-D1, most existing works focus on animating 3DGS while neglecting the physical attributes of the Gaussian primitives themselves (Sec. V-A). Some pioneering works attempt to introduce velocity attributes [178] and Newtonian dynamics rules [191], but this is not sufficient to fully describe the physical motion of 3DGS in space. A potential solution is to introduce more physical attributes in Gaussian primitives, such as material [216], acceleration, and force distribution, which can be regularized by priors from certain simulation tools and physics knowledge. Physically-based rendering is also a direction worth attention, as it enables 3DGS to handle relighting and material editing, producing outstanding inverse rendering results [215]. Future works can explore decoupling geometry and appearance in 3DGS, conducting research from the perspectives of normal reconstruction and the modeling of illumination and materials [216, 273, 91].
VIII-E Realness and Efficiency
Challenges. Realness and efficiency challenges are fundamental issues. They are investigated in various works and have been discussed in Sec. III. In this part, we discuss downstream tasks and techniques optimized for performance and efficiency.
@@ VIII-E Realness and Efficiency
Opportunities. The difficulty in reconstructing clear surfaces has always been a significant challenge affecting rendering realism. As discussed in Sec. V-B, some works have addressed it by attempting to represent surfaces with planar Gaussians. However, this can result in a decline in rendering performance, possibly due to the reduced representational capacity of planar Gaussian primitives or the training instability. Therefore, designing Gaussian primitives better suited for surface representation and introducing a multi-stage training paradigm along with regularization are potential solutions. Storage efficiency is one of the critical bottlenecks of 3DGS. Existing works focus on introducing VQ techniques and compressing SH parameters, as discussed in Sec. III-A1. However, such approaches inevitably affect performance. Therefore, exploring how to design more efficient representations based on 3DGS is a potential way to enhance efficiency [16, 17] while maintaining performance.
VIII-F Other Technical Challenges and Opportunities
@@
An increasing number of research and engineering projects have found that initialization is important in 3DGS. Traditional SfM is not suitable for constrained scenarios, such as sparse view settings, AIGC, and low-light reconstruction. Therefore, more robust initialization methods should be designed to replace original initialization in these scenarios. And Splatting also plays an important role in 3DGS, but it is rarely mentioned in existing works [274, 240]. Designing efficient parallel splatting strategies on pretrained 3DGS has the potential to impact rendering performance and efficiency.
IX Conclusion
@@
The burgeoning interest in the field of 3D Gaussian Splatting (3DGS) has precipitated the emergence of a myriad of related downstream tasks and technologies, thereby contributing to increasing complexity and confusion within the domain. This complexity manifests in various forms, including similar motivations across different works, the incorporation of similar technologies across disparate tasks, and the nuanced differences and interconnections between various technologies. This survey systematically categorizes existing works based on their motivations and critically discusses associated technologies. Our objective is to elucidate common challenges across tasks and technologies, providing a coherent framework for understanding this evolving field. Additionally, we highlight prospective research avenues to inspire continued innovation in 3DGS.